### v1

# TactileNet: Bridging the Accessibility Gap with AI-Generated Tactile Graphics for Individuals with Vision Impairment

Abstract
Tactile graphics are essential for providing access to visual information for the 43 million people globally living with vision loss, as estimated by global prevalence data. However, traditional methods for creating these tactile graphics are labor-intensive and struggle to meet demand. We introduce TactileNet, the first comprehensive dataset and AI-driven framework for generating tactile graphics using text-to-image Stable Diffusion (SD) models. By integrating Low-Rank Adaptation (LoRA) and DreamBooth, our method fine-tunes SD models to produce high-fidelity, guideline-compliant tactile graphics while reducing computational costs. Evaluations involving tactile experts show that generated graphics achieve 92.86% adherence to tactile standards and 100% alignment with natural images in posture and features. Our framework also demonstrates scalability, generating 32,000 images (7,050 filtered for quality) across 66 classes, with prompt editing enabling customizable outputs (e.g., adding/removing details). Our work empowers designers to focus on refinement, significantly accelerating accessibility efforts. It underscores the transformative potential of AI for social good, offering a scalable solution to bridge the accessibility gap in education and beyond.
Figure 1: Examples of our image-to-image translation framework: Top row shows reference natural images, middle row displays TactileNet’s sourced/benchmark tactile graphics, and the bottom row presents generated tactile graphics using our adapters.
1 Introduction
Ensuring visual accessibility for individuals with visual impairments is an increasingly critical challenge in the digital age. According to the International Agency for the Prevention of Blindness (IAPB), an estimated 1.1 billion people worldwide were living with vision loss in 2020, a number projected to increase by 55% by 2050 due to population growth and aging IAPB (2020). Among these, 43 million people are blind, and 295 million have moderate to severe visual impairment, underscoring the urgent need for inclusive solutions. Many individuals with vision loss face significant barriers in education and daily life, as learning materials and information systems remain heavily reliant on visual content. This growing disparity highlights the necessity for scalable, innovative approaches to improve accessibility and bridge the gap for those with visual impairments.
Tactile graphics, which convey visual information through textured surfaces, play a crucial role in bridging this accessibility gap. To be effective, they must adhere to Braille Authority of North America (BANA) guidelines BANA (2025) to ensure clarity and usability. However, traditional production methods remain labor-intensive, time-consuming, and difficult to scale. While modern advancements, such as graphic design software CorelDRAW (2025); AdobeIllustrator (2025); TactileView (2025) and electronic embossers IndexBraille (2025); ViewPlus (2025), have improved efficiency, they fall short in meeting the growing demand for high-quality tactile materials. Automated solutions, including AI-driven approaches and refreshable tactile displays, offer promise but face challenges related to high costs and limited training datasets Mukhiddinov and Kim (2021). Consequently, these technologies have yet to be widely integrated into educational systems.
To address these scalability challenges, we introduce TactileNet, the first comprehensive digital dataset designed to train adapters for AI models that generate tactile images from text prompts or combined text and natural image inputs. Given the absence of paired datasets, these models are initially trained using text-to-image Stable Diffusion (SD) models Ho et al. (2020); Rombach et al. (2021) and later evaluated under image-to-image translation scenarios to ensure practical usability. This step is crucial for preserving essential structural information while omitting extraneous details like color and complex textures, which may hinder tactile perception.
A key innovation of our approach is the integration of Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), which enhance model efficiency and precision for high-fidelity tactile graphic generation. LoRA reduces computational costs by adapting low-rank parameters, while Dreambooth enables personalized training with minimal data. The model’s ability to transform RGB images into tactile formats is validated through human expert feedback, demonstrating its potential to improve tactile graphic quality and accessibility. By automating and streamlining the creation process, our method allows designers to focus on refinement and customization, significantly reducing time and effort.
The contributions of this paper are threefold:
  • Introducing TactileNet, a consolidated dataset for training AI models tailored for tactile applications.
  • Developing a novel methodology leveraging SD, LoRA, and Dreambooth to automate tactile graphic creation and editing through text, aiding tactile graphic designers.
  • Implementing a web-based evaluation protocol to assess model effectiveness through expert human feedback.
2 Related Work
2.1 Tactile Graphics: Current Approaches and Limitations
Recent advancements in tactile graphics production have significantly improved accessibility for blind and visually impaired (BVI) individuals. A key focus has been on ensuring the quality of tactile graphics, with methods like those proposed by Gonzalez et al. (2019) assessing image complexity to determine suitability for conversion. However, these approaches often rely on manual or semi-automated processes, which are labor-intensive and difficult to scale.
The integration of AI into accessibility tools has introduced innovative solutions, such as touchscreen tablets with audio feedback Guinness et al. (2019) and tactile graphics finders Felipe and Guerra-Gómez (2020). These tools enhance the comprehension of visual information through alternative modalities, yet they remain limited in addressing the full pipeline of tactile graphic production. For instance, video description tools Yuksel et al. (2020a, b) provide valuable context but do not automate the creation of tactile graphics.
2.2 GenAI and Deep Learning in Accessibility
Generative AI (GenAI), particularly deep learning (DL), has shown promise in automating complex tasks across various domains Kirillov et al. (2023); Dastjerdi et al. (2024); Khan et al. (2024). However, a major challenge in developing AI models for tactile image generation is the scarcity of high-quality paired datasets. The systematic review by Mukhiddinov and Kim (2021) highlights the limited availability of large-scale datasets and the high cost of refreshable tactile displays, both of which hinder advancements in AI-driven tactile graphics generation.
Text-to-image and image-to-image models, such as Stable Diffusion (SD) Ho et al. (2020); Rombach et al. (2021), have demonstrated the ability to generate high-quality visual content from textual or visual inputs. However, their application in tactile graphics remains underexplored, primarily due to the lack of specialized datasets and the unique structural constraints of tactile representation.
2.3 Gaps in Existing Solutions and Our Contribution
While existing tools address specific aspects of BVI accessibility, they lack a comprehensive solution for scalable, high-quality tactile graphic production. Our work bridges this gap by introducing a curated dataset designed to train AI models for automating tactile graphic creation. Unlike previous approaches, our method leverages SD models, enhanced with Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), to ensure efficiency and precision. This approach not only reduces the time and effort required by tactile graphic designers but also establishes a new benchmark for inclusivity and accessibility in the field.
3 Methodology
In this section, we introduce the preliminaries and outline our methodology for curating the TactileNet dataset and adapting Stable Diffusion (SD) models to generate tactile graphics.
3.1 Preliminaries
3.1.1 Denoising Diffusion Probabilistic Models (DDPMs)
A DDPM Sohl-Dickstein et al. (2015); Ho et al. (2020) consists of two Markov chains: a forward chain that gradually adds noise to data, transforming it into a Gaussian distribution, and a reverse chain that learns to denoise and reconstruct the data Koller and Friedman (2009).
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (1)
Here, $\mathbf{x}_{0}$ represents the original data, $\bar{\alpha}_{t}$ controls the noise level at step $t$, and $\boldsymbol{\epsilon}$ is Gaussian noise. The reverse process, also referred to as denoising, reconstructs the original data by progressively estimating and removing the added noise:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}*\epsilon_{\theta}(x_{t},t)\right)$ (2)
In this equation, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the noise component predicted by the neural network model.
3.1.2 Text-to-Image SD Models
Text-to-Image SD models extend DDPMs, using textual prompts to guide image generation and enable text-to-visual synthesis. These models represent a convergence of vision-language technologies, where the structured approach of DDPMs is adapted to understand and generate images based on textual cues Radford et al. (2021). This innovative application involves not only the transformation of noise into structured images but also the integration of linguistic elements to ensure that the generated visuals accurately reflect the described scenarios Yang et al. (2024); Kawar et al. (2023); Du et al. (2022). The progression from data to noise and back to refined images exemplifies the adaptive use of diffusion techniques in bridging textual and visual modalities.
3.1.3 Fine-Tuning SD Models
Fine-tuning adapts pre-trained models to specific tasks, enhancing performance in applications requiring precision and detail. Our framework leverages two methods: Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), fine-tuned using Kohya’s Trainer Linaqruf (2023) on text-image pairs to refine the model’s ability to interpret textual prompts. During inference, the system supports both text-to-image and image-to-image generation (e.g., via the SD web interface AUTOMATIC1111 (2025)), enabling tactile graphic synthesis from textual descriptions or reference images. This flexibility ensures practical usability, bridging modalities for tactile design.
3.1.4 Low-Rank Adaptation (LoRA)
LoRA is a parameter-efficient fine-tuning method, ideal for resource-constrained scenarios or small datasets. LoRA updates only a subset of the model’s parameters, enabling adaptation to new tasks or improved performance on specific data without comprehensive re-training. Initially, a pre-trained model, which has been trained on a large dataset, serves as the basis for adaptation. LoRA specifically targets the weight matrices in the attention and feed-forward layers of the model layers that are critical for performance. Instead of altering the original weight matrices directly, LoRA introduces a pair of low-rank matrices $A$ and $B$ for each weight matrix $W$ that needs adaptation. For a weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll m,n$. The adapted weight matrix becomes $W+\Delta W$, where $\Delta W=AB$. The low-rank approach ensures that only a fraction of the original parameters are updated, thus significantly reducing the number of trainable parameters. The original parameters of the pre-trained model remain frozen, meaning they are not updated during the adaptation process. Only the parameters of the low-rank matrices $A$ and $B$ (and possibly biases) are trainable, which requires considerably fewer computational resources compared to full model re-training, while maintaining high accuracy.
3.1.5 DreamBooth
DreamBooth fine-tunes SD models using a small set of subject images, enabling the generation of customized images with high fidelity. It involves selecting a unique identifier for a specific subject, which conditions the model to associate the subject’s detailed characteristics and context with this identifier. During the fine-tuning process, the model is trained on a relatively small dataset featuring images of the subject. This targeted training adjusts the model’s parameters, enabling it to generalize the appearance and context of the new subject across various scenarios not directly presented in the training data. This method effectively integrates the subject into the model’s latent space, allowing for the generation of new, high-quality images of the subject in diverse contexts and settings upon request.
Both LoRA and DreamBooth tailor generative models to specific applications. LoRA is ideal for situations requiring limited resources or minor adaptations, affecting only a minimal number of parameters. DreamBooth, although more computationally intensive, enables deep integration of new subjects into the model, facilitating the creation of customized, contextually precise images from text descriptions.
3.2 Dataset Creation
Figure 2: Flow diagram illustrating the process of data compilation from initial sourcing to final dataset compilation.
3.2.1 Data Collection
We curated the TactileNet dataset by sourcing tactile images from online digital libraries, including Perkins College for the Blind, the Tactile Graphics Image Library of American Printing House (TGIL APH), the Provincial Resource Centre for the Visually Impaired (PRCVI), and BTactile Perkins (2025); APH (2024); PRCVI (2024); BTactile (2025). These libraries are known for their high-quality, expert-designed tactile images, which are widely used in educational settings for the visually impaired.
Given the limited number of available images for certain classes (e.g., the ”dog” class in TGIL APH yielded only four images, including a circus dog), we expanded our dataset by adding visually similar images from online social platforms such as Pinterest. To ensure these additional images met the high standards necessary for tactile interpretation, each image was carefully compared against the benchmarks set by the tactile images from the aforementioned libraries, considered as the gold standard in tactile design. This approach ensured consistency in quality and educational value across the dataset. Figure 2 illustrates the process from initial image sourcing to final tactile graphics compilation. Through this careful process, we compiled a dataset of 1,029 tactile images across 66 classes to support robust model training. Key statistics, including class distribution, are provided in Table 2.
Collaboration with Industry Partners
Throughout the development of our dataset, we actively collaborated with industry partners specializing in accessibility solutions to ensure the reliability and real-world applicability of our data. These partners included tactile graphics designers and educators with expertise in visual impairments, who provided critical insights into the practical needs of end-users. Their feedback ensured that our dataset not only meets technical standards but also aligns with the requirements of real-world educational settings for the visually impaired. This collaboration underscores our commitment to creating socially impactful AI solutions that address accessibility challenges.
Figure 3: Comprehensive workflow of our framework, starting with fine-tuning (left), where TactileNet data (tactile images, prompts) refine the SD model. The process transitions to the generation phase (right), applying fine-tuned adapters atop the frozen SD model for text-to-image and image-to-image tactile graphic generation.
3.2.2 Text Prompts Generation for Tactile Graphics
To prepare the dataset for fine-tuning, we generated text prompts for each tactile image using ChatGPT OpenAI (2025) integrated with DALL-E, following a structured template mentioned below.
Create a tactile graphic of an [object], specifically designed for individuals with visual impairments. The graphic should feature raised, smooth lines to delineate the [pattern/features], against a simplistic background to ensure stark contrast.
Here the [object] is replaced by the class name (e.g cat). The [patterns/features] are replaced by specific features (e.g whiskers, eyes, paws etc) in the tactile graphic. These prompts are then thoroughly reviewed to ensure they strictly adhere to our predefined template and whether they are faithful to the input tactile graphic, thereby maintaining consistency and relevance throughout the dataset.
3.3 Models Development and Image Generation
3.3.1 Fine-Tuning Individual Models for Each Category
Our dataset, comprising text-tactile image pairs, was used to fine-tune 66 distinct models (one per class) using LoRA and DreamBooth (Figure 3 (left)). The “tactile” identifier ensured alignment with tactile features during training. Implemented via the Kohya Trainer Linaqruf (2023), this process adapts the Stable Diffusion-based Anything V3 model AnythingV3 (2024); Rombach et al. (2021), originally fine-tuned for an anime-style generation. We selected Anything V3 after comparing generation quality across SD v1 and v1.5. The Anything V3 fine-tuned adapter produced superior results when used atop the base SD v1.5 model during the generation phase, as detailed in the following subsection.
3.3.2 Generation of Tactile Images
Following fine-tuning, tactile graphics are generated using the fine-tuned SD v1.5 model, as depicted in Figure 3 (right). Our generation phase operates under two settings: (1) class-specific prompts alone and (2) prompts combined with natural images. The inclusion of natural images significantly improved output quality, with 30% of graphics meeting quality standards compared to 15% for text-only prompts. These results highlight the importance of multimodal inputs for tactile graphic synthesis, as further evaluated in Section 4.3.
4 Experimental Settings
This section outlines the configuration and environmental settings utilized in our experiments.
4.1 Fine-tuning Configuration
We fine-tune the Anything V3 AnythingV3 (2024) across 66 distinct classes. Each class, supported by training images ranging from a minimum of 9 (llama class) to a maximum of 102 (helicopter class), with an average of approximately 16 images per class. We levearge the sourced tactile images along with the prompts to enable the development of class-specific adapters. These lightweight modules specialize in generating tactile graphics for each category and are integrated atop frozen versions of SD v1.5 for image generation.
For fine-tuning, the DreamBooth configuration prioritized prior preservation with a loss weight of 1.0, ensuring the model retains its ability to generate generic images while adapting to tactile-specific features. The LoRA setup used a network module with linear dimensions set to 32 and an alpha parameter of 16, which controls the scaling of low-rank updates. This configuration ensures effective parameter adaptation without requiring pre-loaded weights.
4.1.1 Optimization and Hardware Configuration Parameters
The optimization process employed the AdamW8bit optimizer Kingma and Ba (2017), renowned for its efficiency, particularly with a learning rate set at $1\times 10^{-4}$ for the UNet Ronneberger et al. (2015) and $5\times 10^{-5}$ for the text encoder, operating under a constant rate scheduler without warm-up steps. Our experiments for fine-tuning were facilitated by the NVIDIA Tesla T4 GPUs provided by Google Colab. A training batch size of 6 was used, utilizing mixed precision training at FP16 across a maximum of 20 training epochs.
4.2 Image Generation Configuration
The primary configuration for text-to-image generation involved the DPM++ 2M Karras sampling method, with 20 steps at image dimensions of 512×512 pixels at a CFG scale of 7. For image-to-image translation experiments, configurations were similar, except for an additional denoising strength of 0.9 to better retain the structural essence of the original images while incorporating new tactile features.
For some challenging classes, deviations were necessary to ensure optimal structure and clarity. Objects requiring finer details, such as bee wings, elephant textures, and duck feathers on water, utilized higher denoising strengths (0.96–1.0) and adjusted CFG scales (8–10) to improve contrast and edge clarity. Classes with intricate internal features like basketball, dinosaur T-Rex teeth, and camera lenses benefited from additional lineart-based ControlNet modules to refine their contours. Water-based objects, such as the sailboat and floating duck, incorporated negative prompts to suppress excessive water details, ensuring the main subject remained distinct. For simpler structures like egg, book, and hatchback car, a LoRA-only approach sufficed, requiring no ControlNet assistance. We utilize NVIDIA TITAN V GPUs with 12288 MiB of memory for generating the images.
4.3 Evaluation Protocol
To assess the quality and usability of the generated tactile graphics, we developed a custom evaluation protocol in collaboration with tactile graphics designers. Standard metrics such as Fréchet Inception Distance (FID) Heusel et al. (2017) and Structural Similarity Index (SSIM) Wang et al. (2004) are insufficient for this specialized domain, as they do not account for the unique requirements of tactile graphics, such as line clarity, texture, and adherence to accessibility guidelines BANA (2025). Instead, our protocol prioritizes expert evaluation by our industry partners, ensuring that the generated graphics meet the practical needs of visually impaired users.
4.3.1 Interface Design and Rationale
Our evaluation interface presents evaluators pairs of images:
  • Reference Image: A natural image depicting the subject (e.g., a cat standing).
  • Tactile Image: The generated tactile graphic or a sourced tactile graphic from established libraries (e.g., APH, Perkins).
This side-by-side comparison allows evaluators to assess how well the tactile graphic captures the essential features and posture of the subject while adhering to tactile design principles. Evaluators are blind to the source of the tactile graphic (generated vs. sourced), ensuring an unbiased assessment.
4.3.2 Evaluation Questions and Metrics
Evaluators are asked to answer the following questions for each tactile graphic:
Q1: Natural Features and Posture Alignment.
Determine whether the tactile image accurately reflects the natural features and pose depicted in the reference image.
Example: If the reference image shows a cat standing, but the tactile image depicts a cat sitting, this should be marked as ’No’.
Instructions: Select ’Yes’ if the tactile image aligns well with the reference image in terms of pose and essential features (e.g., body posture, visible organs). Select ’No’ if there are discrepancies.
Q2: Adherence to Tactile Graphics Guidelines.
Assess whether the tactile graphic follows established tactile graphics standards (e.g., BANA guidelines BANA (2025)).
Example: A tactile graphic with overly complex patterns that might confuse tactile reading should be marked as ’No’.
Instructions: Select ’Yes’ if the tactile image adheres to the guidelines. Select ’No’ if it fails to meet these standards.
Q3: Quality Rating of the Tactile Image.
Rate the quality of the tactile graphic based on its utility and adherence to tactile representation principles.
Options:
  • Accept as Is: The tactile image meets all quality standards and requires no modifications.
  • Accept with Minor Edits: The image is generally acceptable but requires minor modifications to enhance clarity or adherence to guidelines.
  • Accept with Major Edits: The image requires significant changes to be useful as a tactile graphic.
  • Reject (Useless): The image does not meet the standards for tactile graphics and cannot be salvaged through edits.
Instructions: Choose the option that best describes the state of the tactile graphic.
Q4: Optional Feedback.
Provide detailed comments or suggestions for improving the tactile graphic.
Instructions: Use this section to highlight specific issues (e.g., line clarity, texture) or suggest modifications.
4.3.3 Natural Image Selection and Prompt Generation
Natural images were carefully selected to match the sourced gold-standard tactile graphics from established libraries. A team of five undergraduate students was trained to find the closest matches using Google search queries such as ”side profile of [class].” Ambiguities were resolved through max voting when multiple candidates were identified. For each class, two test samples were prepared that would go with the reference natural image:
  • Sample 1: A generated tactile graphic produced by our class-specific adapters.
  • Sample 2: A sourced tactile graphic from established libraries or open-source platforms.
The generated tactile graphics were produced using two types of prompts:
  • Original Prompt: The prompt generated during the prompt generation phase (Section 3.2.2).
  • Paraphrased Prompt: A refined version of the original prompt, paraphrased using DeepSeek DeepSeek (2025) with the instruction: ”Paraphrase the given prompt for tactile generation while preserving the ’tactile’ subject and essential features.”
For each prompt type, eight tactile graphics were generated, and the best match was selected based on the evaluation criteria as discussed.
4.3.4 Statistical Analysis Plan
The evaluation results were analyzed using binary metrics (Yes/No) for Q1 and Q2, and categorical metrics (Accept as Is, Accept with Minor Edits, etc.) for Q3. Percentages and averages were computed to compare the performance of generated vs. sourced tactile graphics. Detailed results and analysis are presented in the following section.
5 Results and Discussion
In this section, we present results including TactileNet dataset statistics (Table 2) and outcomes from text-to-image and image-to-image translation tasks.
5.1 Image-to-Image Translation Evaluation
Table 1 summarizes the quality ratings for both sourced and generated tactile images. Figure 1 illustrates some samples from our evaluation: the first row presents the natural reference images, the second row displays the sourced tactile graphics, and the third row showcases the tactile graphics generated using our fine-tuned adapters.
Table 1: Quality Ratings for Generated vs. Sourced Tactile Graphics
Category Generated (%) Sourced (%)
Accept as Is 32.14 35.71
Accept with Minor Edits 39.29 39.29
Accept with Major Edits 28.57 21.43
Reject (Useless) 00.00 3.57
Figure 4: Example of human-induced errors in natural image pairing: (Left) Natural image of a Bactrian camel; (Center) Sourced tactile graphic of a Dromedary camel; (Right) Generated tactile graphic adhering to the reference.
5.1.1 Key Findings
Our evaluation reveals that both sourced and generated tactile graphics achieved 100% alignment with natural images in terms of pose and structural features (Q1), confirming their accuracy in representation. However, adherence to tactile accessibility guidelines (Q2) was slightly higher for sourced graphics (96.43%) compared to generated ones (92.86%), suggesting that while our fine-tuned models perform well, further refinements are needed for optimal accessibility.
Additionally, 28.57% of generated graphics required major edits, primarily due to excessive visual complexity, such as unintended 3D perspectives in objects like Chair and Bed (Figure 1), which can hinder tactile interpretability. A small fraction of sourced tactile images (3.57%) were rejected mostly due to human-induced mismatches in dataset curation, where incorrect natural images were paired with tactile graphics, as seen in the camel species (Figure 4). This analysis highlights the strengths of AI-generated tactile graphics while emphasizing key areas for improvement, particularly in simplifying complex structures and refining dataset sourcing for higher semantic accuracy.
Class (Source, Generated) Counts Class (Source, Generated) Counts Class (Source, Generated) Counts
Airplane (10, 55) Apple (11, 28) Ball (25, 83)
Banana (13, 156) Bat (13, 60) Bed (14, 95)
Bee (13, 61) Beluga Whale (11, 27) Bicycle (11, 117)
Bird (16, 115) Boat (18, 19) Book (12, 73)
Bottle (11, 137) Camel (10, 109) Camera (12, 219)
Car (25, 106) Cat (22, 142) Chair (12, 117)
Clover (10, 23) Crab (10, 321) Cup (15, 380)
Dinosaur (20, 184) Dog (21, 119) Door (13, 299)
Duck (12, 399) Egg (17, 87) Elephant (20, 29)
Fish (30, 130) Flower (13, 72) Fox (13, 163)
Giraffe (12, 12) Glasses (12, 44) Guitar (18, 70)
Hammer (19, 133) Hat (12, 73) Headphones (11, 53)
Helicopter (102, 35) Horse (23, 93) Hut (10, 226)
Iron (10, 110) Jellyfish (9, 21) Lamp (10, 38)
Laptop (11, 153) Leaf (11, 127) Llama (10, 137)
Motorcycle (14, 161) Pencil (19, 85) Penguin (12, 149)
Planet (12, 36) Rabbit (10, 205) Ring (10, 17)
Rocket (23, 84) Satellite (10, 49) School Backpack (12, 24)
Scooty (12, 46) Ship (13, 13) Shirt (21, 152)
Shoe (18, 207) Snowflake (11, 56) Soda Cans (12, 35)
Spoon (10, 25) Teddy Bear (16, 39) Train (18, 202)
Tree (22, 107) Watch (11, 83) Umbrella (10, 25)
Total (1029, 7050) Mean (15.4, 123.5) Median (12, 93) Max (102, 399) Min (9, 12)
Table 2: Overview of the dataset statistics: scrapped tactile images (Source) used for training the models and the output tactile images (Generated) of our text-to-image adapters.
5.2 Text-to-Image Translation Results
We generated a total of 32,000 tactile images across 66 classes, with 7050 images retained after an initial non-expert human filtering. Unlike the image-to-image translation setup, no natural image was provided as input; instead, the model relied solely on textual prompts to generate tactile graphics.
Figure 5: Prompt editing for customization: (Left) T-shirt generated using the base prompt; (Center) Logo removed by adding ”logo” as a negative prompt; (Right) Pocket removed by omitting the keyword ”pocket” from the prompt.
The per-class count is summarized in Table 2. Our fine-tuned adapters demonstrated adaptability through prompt-based edits (Figure 5), enabling modifications such as logo removal or pocket additions. The prompt used for generating images was: Base Prompt: Create a tactile graphic of a t-shirt with a pocket for the visually impaired, highlighting the round neckline, pocket edges, and hem with raised lines. Ensure the pocket is a distinct raised rectangle, allowing users to discover and feel the detail on the shirt’s chest.
6 Conclusion
This study marks a significant advancement in the development of AI-driven accessibility tools through the application of Generative Artificial Intelligence, effectively addressing the critical need for scalable, high-quality tactile graphics. Our contributions are threefold:
  • TactileNet Dataset: A first-of-its-kind dataset of 1,029 expert-curated tactile images across 66 classes, enabling AI model training tailored to tactile design principles.
  • Efficient Fine-Tuning: Integration of LoRA and DreamBooth reduces labor costs while achieving 92.86% adherence to tactile guidelines, rivalling manually sourced graphics.
  • Human-Centric Evaluation: A protocol validated by tactile experts ensures usability, with generated graphics requiring no rejections and matching natural images in alignment 100%.
By automating tactile graphic generation, our method accelerates production and democratizes access to educational materials for visually impaired learners. This work illustrates how AI can drive social good, providing a blueprint for inclusive technology that prioritizes human needs. We encourage further research and collaboration in this area to enhance accessibility and empower visually impaired individuals.